OpenStudy (anonymous):

Evaluate the function for the given value of x. If you can explain how you got the answer that would be great! :D f(x) = -6x2 - 4x - 4 , x = - 2

hartnn (hartnn):

directly put x=-2 f(-2)= -6 (-2)^2 -4(-2) -4 now (-2)^2 = (-2)(-2) =4 4(-2)=-8 so, f(-2) =-6(4)-(-8)-4 =-24+8-4 i think u can solve further

OpenStudy (anonymous):